### v2

# Unveiling the Blind Spots: A Critical Examination of Fairness in Autonomous Driving Systems

Abstract.
Autonomous driving systems have extended the spectrum of Web of Things for intelligent vehicles and have become an important component of the Web ecosystem. Similar to traditional Web-based applications, fairness is an essential aspect for ensuring the high quality of autonomous driving systems, particularly in the context of pedestrian detectors within them. However, there is an absence in the literature of a comprehensive assessment of the fairness of current Deep Learning (DL)-based pedestrian detectors. To fill the gap, we evaluate eight widely-explored DL-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable a thorough fairness evaluation, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues related to age. The undetected proportions for adults are 20.14% lower compared to children. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that the bias may exacerbate for children and females towards low brightness and low contrast.
copyright: nonecopyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Autonomous driving systems are on track to become the predominant mode of transportation in the future (Stocco et al., 2020). However, these systems are susceptible to software bugs (Garcia et al., 2020), which can potentially result in severe injuries or even fatalities for both pedestrians and passengers. The unfortunate incident in 2018 involving an autonomous vehicle from Uber serves as a stark reminder of these risks (cas, 2018). Given the safety-critical nature of autonomous driving systems, they have garnered substantial attention from the software testing community (Zhang et al., 2022).
Extensive research efforts have been devoted to the testing of autonomous driving systems. For example, Tian et al. (Tian et al., 2018) introduced DeepTest, which applies image transformation to simulate potential camera noise in autonomous driving scenarios. Zhang et al. (Zhang et al., 2018) developed DeepRoad, a Generative Adversarial Network (GAN)-based approach that generates test images from real-world driving scenes. Zhou et al. (Zhou et al., 2020) proposed DeepBillboard, a system for generating adversarial billboards to induce potential steering errors in autonomous vehicles.
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under-investigated in the literature. From a Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Galhotra et al., 2017).
Fairness issues in autonomous driving systems, such as a higher accuracy in detecting pedestrians of white ethnicity compared to black ethnicity, can perpetuate discriminatory outcomes and unequal treatment based on race. This can result in harm to individuals belonging to marginalized groups, further exacerbating existing social inequalities. Therefore, it is crucial to prioritize fairness testing in autonomous driving systems.
To fill the gap, we investigate fairness of eight state-of-the-art DL-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022b). To enable this investigation, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
Our study reveals the following findings: (1) Overall, state-of-the-art pedestrian detectors exhibit significant bias across different age and skin tone groups. On the four datasets examined, the undetected proportions for children surpass those for adults by an average of 20.14%. (2) However, the overall performance of these pedestrian detectors in detecting males and females and dark-skin and light-skin groups does not exhibit a large difference, with only a 1.19% and 0.44% gap in undetected proportions. (3) Regarding the exploration of common scenarios, the detection bias towards the children and females increases under low-brightness and low-contrast conditions compared to adults and males. For instance, the difference in undetected proportions for adults and children increases from 22.45% to 25.48% between day time and night time scenarios.
To summarise, we make the following contributions:
  • We conduct the first comprehensive study of fairness in autonomous driving systems across various datasets and demographic attributes, evaluating eight widely-studied DL-based pedestrian detectors and uncovering significant fairness issues.
  • We augment four real-world datasets with manually labeled demographic information, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels.
  • We publicly release the data, demographic labels, and code used in this study (git, 2024) to facilitate future research on fairness of autonomous driving systems.
2. PRELIMINARIES
This study resides at the intersection of two increasingly important SE topics: software fairness and autonomous driving testing. To provide the necessary context, we begin by reviewing the background knowledge and relevant prior research in these areas.
2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020; Zhang et al., 2023; Zhang and Sun, 2022). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US law (Wick et al., 2019). Consequently, prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2022a). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2022a). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring adult groups over child groups. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
Presently, the majority of software fairness testing research primarily revolves around tabular data, as indicated by a recent survey (Chen et al., 2022a). This paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are more recognizable in autonomous driving datasets and have been widely considered in the fairness literature (Soremekun et al., 2022).
2.2. Autonomous Driving Testing
Autonomous driving is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2019; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022). These techniques mainly involve the automatic generation of artificial driving scenarios to simulate real-world conditions, enabling the evaluation of system behavior and the detection of blind spots or corner bugs across various autonomous driving scenarios.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving. Pedestrian detection is a crucial process identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores (Dollár et al., 2012; Zhang et al., 2016). Brandao (Brandao, 2019) discussed fairness in pedestrian detection but focused solely on traditional detection techniques, which are no longer dominant in contemporary practice. Wilson et al. (Wilson et al., 2019) examined skin tone bias but with limitations to one dataset. Similarly, Kogure et al. (Kogure et al., 2022) only explored age bias with one detection method and a small-scale dataset, lacking a comprehensive analysis. To summarise, current fairness research in pedestrian detection of autonomous driving is still in its early stages and faces limitations in terms of diversity in pedestrian detectors, testing datasets, and sensitive attributes considered.
To address this knowledge gap, our paper presents the a systematic empirical study on group fairness testing in pedestrian detection. We conduct experiments using eight popular detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. Through this comprehensive study, we gain insights into the potential bias and risks associated with using existing pedestrian detectors.
3. Experimental Design
This section introduces our research questions and experimental settings.
3.1. Research Questions
We aim to answer the following research questions (RQs).
RQ1 (Overall fairness): To what extent do widely-studied pedestrian detectors exhibit unfairness concerning common sensitive attributes? This RQ explores the performance difference of widely-studied pedestrian detectors when they are applied to different demographic groups, characterized by sensitive attributes including gender, age, and skin tone.
RQ2 (Fairness in different scenarios): What fairness do pedestrian detectors achieve in different brightness, contrast, and weather conditions? We further investigate the fairness of commonly-studied pedestrian detectors in different autonomous driving scenarios by considering a variety of different brightness (RQ2.1), contrast (RQ2.2), and weather conditions (RQ2.3) (Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020).
In the following, we introduce the pedestrian detectors (Section 3.2), benchmark datasets (Section 3.3), evaluation metrics (Section 3.4), and statistical analysis (Section 3.5) that we use to address the above research questions.
3.2. Pedestrian Detection Methods
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories (Hasan et al., 2020): general object detectors and pedestrian-specific detectors. Next, we briefly introduce each category and its pedestrian detectors we study. Table 1 provides an overview of these detectors (“Detector” shows the name of a pedestrian detector; “Backbone/Model” represents the architecture of the detector, which could be either the specific variant of the model employed, such as ‘YOLOX-Tiny’, ‘YOLOX-S’, ‘YOLOX-L’, or ‘YOLOX-X’ in the case of the YOLOX series, or the pre-trained deep neural network used for extracting features from input images or the model size we choose; “Source” indicates the framework/toolkit name and its source for a given pedestrian detector).
Table 1. Pedestrian detectors.
Type Detector Backbone/Model Source
General object detectors YOLOX YOLOX-X MMDetection (MMDetection Developers, nd)
RetinaNet X-101-64x4d-FPN
Faster R-CNN X-101-64x4d-FPN
Cascade R-CNN X-101-64x4d-FPN
Pedestrian -specific detectors ALFNet ResNet50 ALFNet (liuwei16, 2018)
CSP ResNet50 Pedestron (hasanirtiza, 2021)
MGAN VGG16
PRNet ResNet50 PRNet (sxlpris, 2020)
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two genres (Zou et al., 2023): two-stage and one-stage detectors. Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster R-CNN (Ren et al., 2015) (one of the pioneering detectors in the R-CNN family) and Cascade R-CNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
Pedestrian-specific detectors: Pedestrian-specific detectors use additional pedestrian-related information to improve detection performance (Hasan et al., 2020). In this study, we investigate state-of-the-art pedestrian-specific detectors, including ALFNet (Liu et al., 2018), CSP (Liu et al., 2019), MGAN (Pang et al., 2019), and PRNet (Song et al., 2020). ALFNet uses progressive detection heads on SSD (Liu et al., 2016) to refine initial anchors for improved detection accuracy. CSP introduces an anchor-free approach by locating center points and scaling pedestrians. MGAN uses visible-area bounding-box information to guide attention mask generation for occluded pedestrian detection. PRNet presents a novel progressive refinement network for occluded pedestrian detection.
3.3. Benchmark Datasets
3.3.1. Dataset Selection
We perform our experiments on four real-world datasets that have been extensively studied by researchers to evaluate the performance of pedestrian detectors in autonomous driving (Cao et al., 2020; Hasan et al., 2020). These datasets consist of street-level images captured by cameras mounted on autonomous vehicles, showcasing pedestrians in diverse poses, sizes, and occlusion scenarios. Table 2 presents details about these datasets, including the sensitive attributes, the number of images in each dataset, and the respective time of day when these images were captured. Next, we briefly introduce each dataset:
Table 2. Benchmark datasets.
Name Sensitive Attributes #Images Time
CityPersons gender, age 1,525 day
EuroCity-Day gender, age 2,427 day
EuroCity-Night gender, age 2,126 night
BDD100k gender, age, skin 2,233 day, night
  • CityPersons dataset (Zhang et al., 2017) stands as the most widely-studied benchmark for evaluating pedestrian detectors (Cao et al., 2020; Hasan et al., 2020; Liu et al., 2018; Song et al., 2020; Liu et al., 2019; Pang et al., 2019). Its test set includes 1,525 images captured across six cities, showcasing diverse weather conditions and street scenes.
  • EuroCityPersons dataset (Braun et al., 2019) contains 4,553 images gathered from seven European cities, encompassing both day and night time captures. The dataset can be categorized into two sets: 2,427 images captured during the day, and 2,126 images captured at night, referred to as the EuroCity-Day dataset and the EuroCity-Night dataset, respectively.
  • Berkeley Driving dataset (a.k.a., BDD100k dataset) (Yu et al., 2020) is an extensive driving dataset, including 2,233 images from 40 classes that are typical of driving scenes. These images were captured from four different cities, providing various times of the day. Notably, this dataset showcases a greater diversity of pedestrians than the CityPersons and EuroCityPersons datasets, including individuals with varied skin tones.
3.3.2. Sensitive Attribute Labeling
As introduced in Section 2.1, we focus on three sensitive attributes (i.e., gender, age, and skin tone) of pedestrians that are recognizable in autonomous driving images and widely studied in the fairness literature (Soremekun et al., 2022). To enable fairness testing, we need datasets with labels that identify these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset. Therefore, we manually label the other sensitive attributes for each of the datasets considered in this study.
The labeling process involves two annotators to minimize the influence of labeling bias. We focus on images that align with the widely-adopted “reasonable subset” principle (Dollár et al., 2012), meaning that we label images containing labeled pedestrians with a height of at least 50 pixels and little to none occlusion. For such images, human annotators can label the sensitive attributes with high confidence (Brandao, 2019). Using the filtered datasets, the two annotators independently label the gender and age attributes for each image. For gender, we follow previous studies (Brandao, 2019) and consider only two labels: male and female. As for age, in line with the literature (Brandao, 2019; Kogure et al., 2022), we classify pedestrians into two labels: child and adult, based on their physical characteristics depicted in the images.
To ensure the reliability of the labeling procedure, both annotators independently label the age and gender attributes for each image. We use Cohen’s Kappa ($\kappa$) (Cohen, 1960), a widely-adopted metric for measuring inter-rater agreement, during the independent labeling process. The obtained $\kappa$ values for age and gender attributes in each of the four datasets are summarized in Table 3. According to the literature (Landis and Koch, 1977), a $\kappa$ value between 0.81 and 1 signifies almost perfect agreement, while a value between 0.61 and 0.8 indicates substantial agreement. In our labeling process, we achieve substantial agreement in gender labeling for the EuroCity-Day dataset and almost perfect agreement for all other tasks. This high level of agreement underscores the reliability of our labeling procedure (Wen et al., 2021; Chen et al., 2021). In cases where the two annotators encounter conflicts, an arbitrator is involved in the discussion to reach a consensus. After the labeling process, the summary of the number of labeled pedestrian instances for each dataset is presented in Table 4.
Table 3. Cohen’s $\kappa$ values for labeling age and gender.
Dataset $\kappa_{age}$ $\kappa_{gender}$
CityPersons 0.847 0.814
EuroCity-Day 0.925 0.800
EuroCity-Night 0.847 0.870
BDD100k 0.828 0.854
Table 4. Number of labeled pedestrian instances per dataset.
Gender Age Skin
Dataset Male Female Adult Child LS DS
CityPersons 2,357 1,822 4,568 233 - -
EuroCity-Day 1,726 1,646 4,498 100 - -
EuroCity-Night 1,265 1,318 4,165 68 - -
BDD100k 3,457 2,479 6,293 190 2,724 789
Overall 8,805 7,265 19,524 591 2,724 789
3.3.3. Scenario Labeling
To deeply explore the fairness of pedestrian detectors across various driving scenarios (i.e., different brightness, contrast and weather conditions), we also need images labeled with scenario information. We therefore classify the images containing labeled pedestrians (5,933 out of the total 8,311 images in all datasets) into different scenarios, to enable our analysis in RQ2. An alternative approach to achieving this purpose is via generating images with different scenarios using existing test generation techniques from the autonomous driving literature. We do not choose this approach because generated images are not real images and can suffer from unnaturalness (Zhang et al., 2022).
Labeling brightness. Brightness represents the overall lightness or darkness of the image. To distinguish the brightness of the images, we categorize the datasets into day time and night time. As seen in Table 2, images in the CityPersons and EuroCity-Day datasets belong to day time, while images in the EuroCity-Night dataset belong to night time. For the BDD100k dataset, Wilson et al. (Wilson et al., 2019) have publicly provided the labels for day time and night time. We directly use the label information for study.
Labeling contrast. Contrast is the difference in brightness between objects in an image. To quantify the contrast of each image, we use the RMS contrast measurement (Peli, 1990), a standard measure in the computer vision literature (Bhuiyan and Khan, 2018; Jakhetiya et al., 2018; Pelli and Bex, 2013). To apply the RMS measurement, we first need to convert all images into the grayscale mode (Peli, 1990). Then, we calculate the RMS contrast value for each image based on the converted version. A higher RMS contrast value indicates a greater contrast. We partition the range of RMS contrast values into three classes, each class featuring equal value intervals. These classes correspond to distinct levels, labeled from level 1 to level 3, with higher levels indicating images with higher contrast. Then, we categorize the images into their respective contrast levels based on their RMS contrast values.
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2019). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because the two annotators and the arbitrator all cannot accurately distinguish the weather conditions.
To measure inter-rater agreement during manual labeling, we also use Cohen’s Kappa ($\kappa$). The $\kappa$ value is 0.813, indicating almost perfect agreement (Landis and Koch, 1977). This high level of agreement confirms the reliability of our labeling procedure. After scenario labeling, the summary of the number of images under different brightness, contrast, and weather conditions is presented in Table 5.
Table 5. Number of images in different brightness conditions, contrast levels, and weather conditions.
Brightness conditions Contrast levels Weather conditions
Day-time Night-time Level 1 Level 2 Level 3 Rainy Non-Rainy
4409 1524 1163 3933 837 277 3800
3.4. Evaluation Metric
We employ Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD), which are the most widely-adopted fairness metrics in the literature (Hort et al., 2023). SPD quantifies the difference in probabilities of favorable outcomes between unprivileged and privileged groups; EOD indicates the difference in the true-positive rate between unprivileged and privileged groups. In the context of pedestrian detection, both SPD and EOD measure the disparity in proportions of successfully detected pedestrians between privileged and unprivileged groups. They also both express the difference in miss rates between privileged and unprivileged groups.
Miss Rate (MR) is the most commonly-studied performance metric in pedestrian detection (Brandao, 2019; Kogure et al., 2022), which quantifies the proportion of undetected pedestrians. Formally, it is calculated as follows:
(1) $MR=1-\frac{TP}{TP+FN},$
where TP (true positives) refers to the number of successfully detected ground-truth bounding boxes, and FN (false negatives) denotes the number of undetected ground-truth bounding boxes. Pedestrian detectors generate bounding box locations and confidence scores for recognized “person” instances in images. To assess whether a given ground-truth bounding box is successfully detected, the standard method in the literature is to use the Intersection over Union (IoU) metric (Dollár et al., 2012). The IoU metric quantifies the degree of overlap between the ground-truth bounding box and the detected bounding box. If the IoU value is greater than 50%, the ground-truth bounding box is considered successfully detected (Dollár et al., 2012). Otherwise, it is classified as undetected.
3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups) we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely adopted to analyze differences between two proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if the obtained $p$-value falls below 0.05.
4. Results
This section answers our RQs based on experimental results.
4.1. RQ1: Overall Fairness
RQ1 investigates the overall fairness of eight state-of-the-art pedestrian detectors with respect to gender, age, and skin tone. First, for each detector, we compute the miss rate (MR) for different demographic groups and calculate EOD based on the MR results over all the datasets that we study. We also use the two-proportion z-test to determine the significance of any observed unfairness (as described in Section 3.5). Table 6 presents the results, with significant unfairness results highlighted in shading. Next, we analyze the results for gender, age, and skin tone, respectively.
Table 6. RQ1: Overall fairness with respect to gender, age and skin tone. Statistically significant unfairness (i.e., MR difference) is highlighted in shading.
Detectors MR Male MR Female EOD (Gender) MR Adult MR Child EOD (Age) MR Light-skin MR Dark-skin EOD (Skin tone)
YOLOX 9.78% 10.50% -0.72% 12.64% 42.47% -29.83% 5.21% 3.80% 1.41%
RetinaNet 10.72% 12.59% -1.87% 14.36% 44.33% -29.97% 8.33% 4.44% 3.90%
Faster RCNN 3.80% 4.13% -0.32% 5.24% 26.06% -20.82% 5.91% 3.30% 2.62%
Cascade RCNN 3.87% 4.16% -0.28% 5.12% 26.57% -21.44% 5.21% 3.04% 2.17%
ALFNet 30.86% 32.90% -2.04% 36.62% 53.47% -16.85% 42.55% 43.98% -1.43%
CSP 33.49% 35.25% -1.76% 37.74% 50.42% -12.68% 61.38% 64.77% -3.39%
MGAN 29.76% 30.97% -1.21% 33.58% 46.53% -12.95% 52.94% 54.88% -1.94%
PRNet 40.28% 41.62% -1.34% 44.69% 61.25% -16.56% 59.65% 59.44% 0.21%
Average 20.32% 21.52% -1.19% 23.75% 43.89% -20.14% 30.15% 29.71% 0.44%
[…]
Then, we explore the difference across the four datasets used in our study. We find that in the CityPersons and EuroCity-Day datasets, which exclusively contain day-time data, there are no significant differences in the miss rate between females and males. However, the results on the EuroCity-Night dataset present a contrasting observation, where all detectors demonstrate a higher miss rate for females, thus illustrating bias in detecting females. In the BDD100k dataset that includes both day-time and night-time images, the performance difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
[…]
Furthermore, we illustrate the miss rate difference across the four datasets that we use in Figure 1, with significant age unfairness indicated by labeled miss rate values. From the figure, we observe that the miss rate of children is consistently higher than that of adults for all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant difference in the miss rate with $p$-value ¡ 0.05. This indicates that the bias favoring adults is not specific to a particular dataset or model, highlighting a strong unfairness between adults and children.
Figure 1. RQ1: Miss rates for children and adults in each dataset. Statistically significant age unfairness (i.e., miss rate difference) is indicated by labeled miss rate values.
The age bias can be attributed to the inherent challenge of detecting small objects in object and person detection research, owing to the limited information provided by small bounding box sizes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To verify this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 2, reveal a correspondence between the two distributions. Children as well as undetected pedestrians tend to have smaller bounding boxes.
Figure 2. Bounding boxes distribution of adults and children (left) and bounding boxes distribution of undetected and detected pedestrians (right).
Skin tone. As shown in Table 6, current pedestrian detectors indicate an impartial performance between light-skin and dark-skin individuals, with a minor average miss rate discrepancy of 0.44% between light-skin and dark-skin groups. Within the scope of eight detectors assessed, three were observed to have marginally increased miss rates for the light-skin group relative to the dark-skin group (with three of them being statistically significant).
Ans. to RQ1: On average, the pedestrian detectors currently in use do not show significant performance differences when considering gender and skin tone. However, a pronounced disparity is observed across different age groups. Specifically, the data points to a substantial miss rate of 20.14% when detecting between children and adults. This significant margin underscores a heightened risk to child pedestrians within the ambit of present autonomous driving systems.
4.2. RQ2: Fairness in Different Scenarios
RQ2 investigates the fairness of state-of-the-art pedestrian detectors under different real-world autonomous driving scenarios.
4.2.1. RQ2.1: Different brightness conditions.
As introduced in Section 3.3.3, we consider brightness conditions from two perspectives: day time and night time.
We first evaluate the average miss rates of the eight pedestrian detectors under day time and night time. The results, presented in Table 7, indicate a noticeable increase in average miss rates during night time compared to day time. For example, the average miss rates for males and females at night are 33.20% and 36.57%, respectively, compared to 17.24% and 17.32% during the day. A similar pattern is observed for age and skin tone. This indicates that the transition from day time to night time influences the performance of pedestrian detectors, with statistically significant higher miss rates observed at night.
Table 7. RQ2.1: MR and EOD of each pedestrian detector under day time and night time. The statistically significant unfairness results are highlighted in shading.
Detectors Gender Age Skin Tone (LS: Light Skin, DS: Dark Skin)
Day-time Night-time Day-time Night-time Day-time Night-time
MR Male MR Female EOD MR Male MR Female EOD MR Adult MR Child EOD MR Adult MR Child EOD MR LS MR DS EOD MR LS MR DS EOD
YOLOX 9.73% 9.38% 0.35% 9.99% 14.53% -4.54% 10.80% 40.77% -29.97% 17.81% 54.93% -37.12% 4.89% 4.03% 0.86% 7.06% 2.52% 4.53%
RetinaNet 10.42% 11.79% -1.37% 11.99% 15.48% -3.49% 12.73% 41.92% -29.19% 18.97% 61.97% -43.01% 7.44% 4.33% 3.11% 13.38% 5.04% 8.34%
Faster RCNN 3.66% 3.59% 0.07% 4.41% 6.06% -1.66% 4.50% 25.00% -20.50% 7.34% 33.80% -26.46% 5.36% 3.13% 2.23% 9.00% 4.20% 4.80%
Cascade RCNN 3.58% 3.68% -0.10% 5.11% 5.87% -0.76% 4.52% 25.58% -21.05% 6.81% 33.80% -26.99% 4.84% 2.99% 1.86% 7.30% 3.36% 3.94%
ALFNet 24.52% 24.39% 0.13% 57.29% 63.42% -6.14% 26.52% 49.42% -22.90% 65.10% 83.10% -18.00% 37.05% 38.66% -1.61% 73.48% 73.95% -0.47%
CSP 26.76% 26.38% 0.38% 61.57% 67.09% -5.51% 28.64% 46.92% -18.28% 63.42% 76.06% -12.64% 55.90% 59.85% -3.95% 92.21% 92.44% -0.22%
MGAN 25.48% 26.29% -0.81% 47.59% 47.76% -0.17% 27.87% 43.46% -15.59% 49.72% 69.01% -19.30% 47.12% 49.70% -2.58% 85.64% 84.03% 1.61%
PRNet 33.73% 33.07% 0.66% 67.63% 72.33% -4.70% 34.80% 56.92% -22.13% 72.60% 92.96% -20.36% 54.39% 54.63% -0.24% 89.29% 86.55% 2.74%
Average 17.24% 17.32% -0.09% 33.20% 36.57% -3.37% 18.80% 41.25% -22.45% 37.72% 63.20% -25.48% 27.31% 27.16% 0.15% 47.17% 44.01% 3.16%
Then, we investigate whether the performance is equally decreased for different demographic groups. Specifically, we explore the fairness change from day time to night time. Table 7 shows the results, with statistically significant unfairness (i.e., miss rate difference) emphasized in shading.
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the night-time condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences.
For age, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night with statistically significant. This suggests a higher probability of children being undetected during night time conditions.
For skin tone, the miss rate difference between dark-skin and light-skin groups increases from day time to night time, with the average difference increasing from 0.15% at day time to 3.16% at night time. This implies a small increased detection difficulty for the light-skin group at night.
Ans. to RQ2.1: Lower brightness conditions amplify the bias towards children and female in pedestrian detection.
4.2.2. RQ2.2: Different contrast levels.
Figure 3. RQ2.2: Average MR of eight pedestrian detectors and EOD trends under different contrast levels.
Following the road map of RQ2.1, we first compare the average miss rates of all pedestrian detectors at different contrast levels, as shown in Figure 3. For skin tone and age, we observe that the average miss rate increases at a lower contrast level (level 1) compared to the other two levels. However, for gender, the average miss rate of all pedestrian detectors does not show an obvious variation across different contrast levels.
Then, we investigate the fairness changes, as indicated by the EOD trends in Figure 3.
For skin tone, as the contrast level decreases, we observe varied responses from pedestrian detectors. Specifically, models like CSP and PRNet exhibit increased difficulty in detecting dark-skin individuals. Conversely, detectors such as RetinaNet and Faster RCNN demonstrate more challenges in identifying the light-skin group. These results demonstrate that contrast conditions have a potential impact on the detection of different skin tone groups
For gender, a reverse trend in EOD emerges. With decreasing contrast levels, pedestrian detectors progressively show a higher miss rate for females than for males. This observation indicates that low contrast conditions negatively impact the detection of females, thereby contributing another factor to the bias and unfairness in pedestrian detection.
For age, the EOD between children and adults maintains a large difference across all contrast levels. Moreover, general object detectors, including YOLOX, RetinaNet, Faster RCNN, and Cascade RCNN, exhibit increased miss rates difference at lower contrast levels. This observation implies that contrast levels in general object detectors could potentially exacerbate the bias, amplifying the disparity in detection performance between adults and children.
Ans. to RQ2.2: Low-contrast conditions have a negative influence on the fairness of some existing pedestrian detectors regarding gender, skin tone and age, especially posing a greater challenge in detecting children and females.
4.2.3. RQ2.3: Different weather conditions.
Table 8 presents the miss rates of the eight pedestrian detectors under non-rainy and rainy weather conditions, with statistically significant unfairness results shaded.
Table 8. RQ2.3: MR and EOD of each pedestrian detector under non-rainy and rainy weather conditions. The statistically significant unfairness results are shaded.
Detectors Gender Age Skin Tone (LS: Light Skin, DS: Dark Skin)
Non-rainy weather Rainy weather Non-rainy weather Rainy weather Non-rainy weather Rainy weather
MR Male MR Female EOD MR Male MR Female EOD MR Adult MR Child EOD MR Adult MR Child EOD MR LS MR DS EOD MR LS MR DS EOD
YOLOX 9.58% 8.95% 0.63% 8.33% 5.17% 3.16% 10.45% 39.95% -29.50% 9.40% 30.00% -20.60% 4.19% 3.47% 0.72% 16.94% 13.51% 3.42%
RetinaNet 10.62% 11.35% -0.74% 10.09% 8.05% 2.04% 12.49% 40.65% -28.16% 13.61% 40.00% -26.39% 6.56% 3.81% 2.74% 21.77% 13.51% 8.26%
Faster RCNN 3.68% 3.29% 0.39% 4.39% 2.87% 1.51% 4.36% 23.09% -18.73% 4.86% 30.00% -25.14% 4.59% 2.60% 1.99% 15.32% 10.81% 4.51%
Cascade RCNN 3.68% 3.48% 0.20% 4.39% 2.87% 1.51% 4.53% 23.79% -19.25% 4.38% 25.00% -20.62% 3.93% 2.60% 1.33% 16.13% 8.11% 8.02%
ALFNet 23.52% 22.92% 0.60% 42.54% 40.80% 1.74% 25.00% 47.34% -22.34% 45.06% 75.00% -29.94% 34.90% 36.92% -2.02% 66.13% 64.86% 1.26%
CSP 25.29% 24.35% 0.94% 50.44% 44.25% 6.19% 26.86% 44.11% -17.25% 48.78% 65.00% -16.22% 53.05% 57.02% -3.97% 79.03% 89.19% -10.16%
MGAN 24.63% 25.26% -0.62% 50.88% 50.57% 0.30% 26.49% 41.11% -14.62% 53.00% 65.00% -12.00% 44.68% 47.83% -3.15% 72.58% 78.38% -5.80%
PRNet 32.73% 31.64% 1.09% 51.75% 50.57% 1.18% 33.19% 53.58% -20.39% 53.32% 75.00% -21.68% 51.99% 52.86% -0.87% 77.42% 81.08% -3.66%
Overall 16.72% 16.41% 0.31% 27.85% 25.65% 2.20% 17.92% 39.20% -21.28% 29.05% 50.63% -21.57% 25.49% 25.89% -0.40% 45.67% 44.93% 0.73%
In terms of the overall detection performance of all pedestrian detectors, the miss rate for each demographic group increases in the rainy weather condition, possibly due to droplets covering the camera and disrupting the detectors. Regarding the fairness, we find that the miss rate difference increases between the children and adults (with EOD from -21.28% to -21.57%) and between males and females (with EOD from 0.31% to 2.20%), indicating a higher miss rate for the children and males from non-rainy to rainy conditions. However, for these increases, we just observe a statistically significant difference from children and adults, the other two attributes do not exhibit statistically significant difference. Therefore, it can be concluded that rainy weather conditions have a minor impact on the fairness of pedestrian detectors.
Ans. to RQ2.3: Rainy weather conditions appear to have a slight effect on the equitable functioning of pedestrian detectors. There is an observed uptick in the discrepancy of miss rates when comparing demographics such as children versus adults, males versus females, and dark-skinned groups versus light-skinned groups. However, these increases are relatively minimal.
5. Discussion
5.1. Fairness-performance Trade-off
Many studies (Ge et al., 2022; Wick et al., 2019; Chen et al., 2023; Hort et al., 2021) have revealed that fairness improvement usually comes at the cost of machine learning performance (e.g., accuracy). Therefore, developers need to grapple with the challenge of optimizing ML performance without compromising fairness, encapsulating this tension as the “fairness-performance trade-off.”
Nonetheless, we have noticed that pedestrian detectors exhibiting greater fairness (i.e., lower EOD values) can achieve superior performance (i.e., lower overall miss rates) under some environmental situations. For example, in the results of RQ2.1, we find that during daytime compared to nighttime, the eight pedestrian detectors achieved better detection results, concurrently displaying reduced EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2, depicted in Figure 3, reveal that heightened contrast levels may lead to improved detection performance and fairness.
Our findings provide a positive contrast to the previous fairness-performance trade-off theory, implicating that “You CAN have your cake and eat it” in the context of pedestrian detection. The immediately actionable insight is the existence of an optimal balance for enhancing both fairness and performance: by adjusting the brightness and contrast conditions of the captured images.
5.2. Implications
Implication for researchers. The significant unfairness exhibited by current pedestrian detectors offers valuable research opportunities for researchers. 1) Fairness improvement via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific image editing techniques to adjust contrast and brightness of input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Fairness improvement via multi-objective optimization. Given our discovery that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. 3) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 4) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage fellow researchers to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
Implication for developers. Fairness is a critical non-functional requirement for Web-based applications, but our study demonstrates the existence of significant bias in state-of-the-art pedestrian detectors. It is imperative for developers to prioritize their efforts in this domain. Addressing these concerns goes beyond enhancing the quality of autonomous driving systems; it also serves to safeguard autnomous drving companies from ethical, reputational, financial, and legal repercussions that may arise in the event of violations of anti-discrimination laws.
Implication for policy makers. Governments have a role in raising awareness about the potential biases in autonomous driving systems. 1) Regulatory measures. Autonomous driving systems play a crucial role in ensuring human safety. Our results reveal that fairness issues exist in modern pedestrian detection models. Therefore, it is essential for policy makers to enact regulations and standards that safeguard the rights of all individuals and address these concerns appropriately. 2) Enhanced protection for vulnerable pedestrians. Policies should also aim to address safety issues for vulnerable pedestrians such as children, who are, as revealed by this study, more likely to be overlooked by the current detection models.
6. Conclusion
This paper presents the first comprehensive study on fairness of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors with respect to gender, age, and skin tone. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. However, we observe balanced detection performance for male and female, light-skin and dark-skin individuals. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our investigation demonstrates that the detection performance for children and females is further reduced under low-brightness and low-contrast conditions, exacerbating the unfairness related to age and gender. Overall, this study sheds light on the fairness issues faced by existing pedestrian detectors, emphasizing the importance of addressing bias related to age and skin tone. The insights gained can pave the way for more fair and unbiased autonomous driving systems in the future.