### v3

# Bias Behind the Wheel: Fairness Analysis of Autonomous Driving Systems

Abstract.
This paper analyzes fairness in automated pedestrian detection, a crucial but under-explored issue in autonomous driving systems. We evaluate eight state-of-the-art deep learning-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable thorough fairness testing, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues, particularly related to age. The undetected proportions for children are 20.14% higher compared to adults. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Moreover, we observe that pedestrian detectors can demonstrate both enhanced fairness and superior performance under specific driving conditions, which challenges the fairness-performance trade-off theory widely acknowledged in the fairness literature. We publicly release the code, data, and results to support future research on fairness in autonomous driving.
Fairness, pedestrian detection, autonomous driving
copyright: acmcopyrightjournal: TOSEMjournalyear: 2024journalvolume: 1journalnumber: 1article: 1publicationmonth: 7price: 15.00price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software creation and managementccs: Computing methodologies Machine learning
1. Introduction
Autonomous driving systems are on track to become the predominant mode of transportation in the future (Stocco et al., 2020). However, these systems are susceptible to software bugs (Garcia et al., 2020), which can potentially result in severe injuries or even fatalities for both pedestrians and passengers. The unfortunate incident in 2018 involving an autonomous vehicle from Uber serves as a stark reminder of these risks (cas, 2018). Given the safety-critical nature of autonomous driving systems, they have garnered substantial attention from the software testing community (Zhang et al., 2022).
Extensive research efforts have been devoted to the testing of autonomous driving systems. For example, Tian et al. (Tian et al., 2018) introduced DeepTest, which applies image transformation to simulate potential camera noise in autonomous driving scenarios. Zhang et al. (Zhang et al., 2018) developed DeepRoad, a Generative Adversarial Network (GAN)-based approach that generates test images from real-world driving scenes. Zhou et al. (Zhou et al., 2020) proposed DeepBillboard, a system for generating adversarial billboards to induce potential steering errors in autonomous vehicles.
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under investigation in the literature. From a Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Chen et al., 2024).
Fairness issues in autonomous driving systems, such as a higher accuracy in detecting male pedestrians compared to females, can perpetuate discriminatory outcomes and unequal treatment based on gender. This can result in harm to individuals belonging to marginalized groups, further exacerbating existing social inequalities. Therefore, it is crucial to prioritize fairness testing in autonomous driving systems.
To fill the gap, we investigate fairness of eight state-of-the-art Deep Learning (DL)-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022). To enable this investigation, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
Our study reveals the following findings: (1) Overall, state-of-the-art pedestrian detectors exhibit significant bias regarding age. On the four datasets examined, the undetected proportions for children surpass those for adults by an average of 20.14%. However, the overall performance of these pedestrian detectors in detecting males and females and dark-skin and light-skin groups does not exhibit a large difference, with only a 1.19% and 0.44% gap in undetected proportions. (2) The studied pedestrian detectors reveal significant gender biases during night time, with a higher proportion of females going undetected compared to males. This situation may aggravate existing societal concerns about female safety during nighttime outings. (3) In contrast to the commonly accepted fairness-performance trade-off, our findings suggest that pedestrian detectors can achieve enhanced fairness and detection performance under specific driving scenarios, such as those with higher brightness levels.
To summarise, we make the following contributions:
  • We conduct the first comprehensive study of fairness in autonomous driving systems across various datasets and demographic attributes, evaluating eight widely-studied DL-based pedestrian detectors and uncovering significant fairness issues.
  • We augment four real-world datasets with manually labeled demographic information, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels.
  • We publicly release the data, demographic labels, and code used in this study (git, 2024) to facilitate future research on fairness of autonomous driving systems.
2. Background and Related Work
This study resides at the intersection of two increasingly important SE topics: software fairness and autonomous driving testing. To provide the necessary context, we begin by reviewing the background knowledge and relevant prior research in these areas.
2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020a; Zhang et al., 2023; Zhang and Sun, 2022; Nguyen et al., 2023). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US anti-discrimination law (Biddle, 2005; Feldman et al., 2015; Dwork et al., 2012; Zafar et al., 2017; Wick et al., 2019). Consequently, testing and prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2024; Wan et al., 2023). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2024). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring adult groups over child groups. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
Recently, Chen et al. (Chen et al., 2024) has presented a comprehensive survey of fairness testing research and analyzed its trend. This survey points out that the majority of existing work revolves around tabular data (Biswas and Rajan, 2020b, 2021; Gohar et al., 2023; Zhang and Harman, 2021; Chen et al., 2023). For example, Biswas and Rajan (Biswas and Rajan, 2020b) evaluated fairness of machine learning models on a crowd sourced platform using tabular datasets covering tasks such as credit risk prediction, income prediction, marketing, and loan application. Similarly, Chen et al. (Chen et al., 2023) conducted an empirical study on the group fairness achieved by state-of-the-art bias mitigation methods across eight commonly used tabular data-driven decision tasks. In the contrast, our paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are recognizable in autonomous driving datasets. These sensitive attributes have been demonstrated to be the most widely considered ones in the fairness testing literature (Chen et al., 2024).
2.2. Autonomous Driving Testing
Autonomous driving is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2017; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022; Zuo et al., 2021). For instance, Tian et al. (Tian et al., 2018) proposed DeepTest, a novel technique using image transformations to emulate potential camera disturbances encountered in driving environments. Zhang et al.  (Zhang et al., 2018) presented DeepRoad, employing Generative Adversarial Networks to craft test images derived from actual driving scenarios. Zhou et al.  (Zhou et al., 2020) introduced DeepBillboard, aiming to generate adversarial billboards that could lead to steering mistakes in autonomous vehicles. Guo et al. (Guo et al., 2022) developed LiRTest, marking the inaugural automated testing technique for LiDAR-equipped autonomous driving software.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only a few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving, indicating that this area remains underdeveloped.
Pedestrian detection is a crucial process that identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores  (Dollár et al., 2012; Zhang et al., 2016). Despite its significance, research on the fairness of pedestrian detection within autonomous driving is limited. Brandao (Brandao, 2019) explored fairness in pedestrian detection, concentrating only on classic machine learning-based methods. These classic techniques, reliant on manually-defined features, have been eclipsed by deep learning-based pedestrian detection approaches, now prevalent in the autonomous driving industry. Wilson et al.  (Wilson et al., 2019) focused on skin tone bias, confined their analysis to a single dataset and two general object detectors. Similarly, Kogure et al.  (Kogure et al., 2022) explored only age bias using a small-scale dataset and a detection method that is no longer state-of-the-art.
In summary, current fairness studies in pedestrian detection suffer from a lack of variety in the pedestrian detectors evaluated, the datasets used, and the range of sensitive attributes explored. Furthermore, no previous work has explored how different environmental characteristics (e.g., brightness, contrast, and weather conditions) affect fairness.
To address this knowledge gap, our paper presents a comprehensive empirical study on revealing fairness issues in pedestrian detection. We conduct experiments using eight popular detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. The scale and diversity of our experiments enable us to provide comprehensive insights into the fairness circumstances of the existing pedestrian detectors.
3. Experimental Design
This section introduces our research questions and experimental design.
Figure 1. Overview of our experimental settings.
3.1. Research Questions
We aim to answer the following research questions (RQs).
RQ1 (Overall fairness): To what extent do widely-studied pedestrian detectors exhibit unfairness concerning common sensitive attributes? This RQ explores the performance difference of widely-studied pedestrian detectors when they are applied to different demographic groups, characterized by sensitive attributes including gender, age, and skin tone.
RQ2 (Fairness in different scenarios): What fairness do pedestrian detectors achieve in different brightness, contrast, and weather conditions? We further investigate the fairness of commonly-studied pedestrian detectors in different autonomous driving scenarios by considering a variety of brightness (RQ2.1), contrast (RQ2.2), and weather conditions (RQ2.3) (Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020).
Figure 1 illustrates our experimental settings to answer these RQs. In the following, we introduce the pedestrian detectors, benchmark datasets, evaluation metrics, and experimental details.
3.2. Pedestrian Detectors
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories (Hasan et al., 2020): general object detectors and pedestrian-specific detectors. Next, we briefly introduce each category and the corresponding pedestrian detectors.
Table 1 provides an overview of these detectors. “Detector” shows the name of a pedestrian detector; “Backbone” represents the pre-trained deep neural network used for extracting features from input images; “Source” indicates the framework/toolkit name and its source for a given pedestrian detector.
Table 1. Pedestrian detectors.
Type Detector Backbone Source
General object detectors YOLOX - MMDetection (MMDetection Developers, nd)
RetinaNet X-101-64x4d-FPN
Faster R-CNN X-101-64x4d-FPN
Cascade R-CNN X-101-64x4d-FPN
Pedestrian-specific detectors ALFNet ResNet50 ALFNet (liuwei16, 2018)
CSP ResNet50 Pedestron (hasanirtiza, 2021)
MGAN VGG16
PRNet ResNet50 PRNet (sxlpris, 2020)
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two genres (Zou et al., 2023): two-stage and one-stage detectors. Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016) used in Apollo’s autonomous driving systems (Apo, 2023)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster R-CNN (Ren et al., 2017) (one of the pioneering detectors in the R-CNN family) and Cascade R-CNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
Pedestrian-specific detectors: Pedestrian-specific detectors use additional pedestrian-related information to improve detection performance (Hasan et al., 2020). In this study, we investigate state-of-the-art pedestrian-specific detectors, including ALFNet (Liu et al., 2018), CSP (Liu et al., 2019), MGAN (Pang et al., 2019), and PRNet (Song et al., 2020). ALFNet uses progressive detection heads on SSD (Liu et al., 2016) to refine initial anchors for improved detection accuracy. CSP introduces an anchor-free approach by locating center points and scaling pedestrians. MGAN uses visible-area bounding-box information to guide attention mask generation for occluded pedestrian detection. PRNet presents a novel progressive refinement network for occluded pedestrian detection.
3.3. Benchmark Datasets
3.3.1. Dataset Selection
We perform our experiments on four real-world datasets that have been extensively studied by researchers to evaluate the performance of pedestrian detectors in autonomous driving (Cao et al., 2020; Hasan et al., 2020). These datasets consist of street-level images captured by cameras mounted on autonomous vehicles, showcasing pedestrians in diverse poses, sizes, and occlusion scenarios. Table 2 presents details about these datasets, including the sensitive attributes, the number of images in each dataset, and the respective time of day when these images were captured. Next, we briefly introduce each dataset:
Table 2. Benchmark datasets.
Name Sensitive Attributes #Images Time
CityPersons gender, age 1,525 day
EuroCity-Day gender, age 2,427 day
EuroCity-Night gender, age 2,126 night
BDD100k gender, age, skin 2,233 day, night
  • CityPersons dataset (Zhang et al., 2017) stands as the most widely-studied benchmark for evaluating pedestrian detectors (Cao et al., 2020; Hasan et al., 2020; Liu et al., 2018; Song et al., 2020; Liu et al., 2019; Pang et al., 2019). Its test set includes 1,525 images captured across six cities, showcasing diverse weather conditions and street scenes.
  • EuroCityPersons dataset (Braun et al., 2019) contains 4,553 images gathered from seven European cities, encompassing both day and night time captures. The dataset can be categorized into two sets: 2,427 images captured during the day, and 2,126 images captured at night, referred to as the EuroCity-Day dataset and the EuroCity-Night dataset, respectively.
  • Berkeley Driving dataset (a.k.a., BDD100k dataset) (Yu et al., 2020) is an extensive driving dataset, including 2,233 images from 40 classes that are typical of driving scenes. These images were captured from four different cities, providing various times of the day. Notably, this dataset showcases a greater diversity of pedestrians than the CityPersons and EuroCityPersons datasets, including individuals with varied skin tones.
3.3.2. Sensitive Attribute Labeling
We focus on three sensitive attribute: gender, age, and skin tone. These attributes are identifiable in autonomous driving images and are recognized as the three most extensively studied sensitive attributes in fairness testing literature (Chen et al., 2024). To enable fairness analysis, we need datasets with labels that indicate these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset.
We manually label gender and age for each of the datasets considered in this study. For skin tone, we only use the BDD100k dataset. As described in Section 3.3.1, this dataset shows a greater diversity in skin tones compared to other datasets. This diversity is linked to the geographic locations where the datasets were collected. Specifically, BDD100k, collected in the United States, exhibits more varied skin tones. In contrast, other datasets from European cities, such as CityPersons collected in Germany, have significantly fewer dark-skin individuals. As reported (pew, 2024; wik, 2024a, b), there is only around 1% representation of individuals with dark skin in Germany. During our labeling process, we indeed encountered scarcity of dark-skin individuals in the CityPersons, EuroCity-Day, and EuroCity-Night datasets, leading to our decision not to label these three datasets with skin tone attributes.
The labeling process involves two annotators to minimize the influence of labeling bias. We focus on images that align with the widely-adopted “reasonable subset” principle (Dollár et al., 2012), meaning that we label images containing labeled pedestrians with a height of at least 50 pixels and little to none occlusion. For such images, human annotators can label the sensitive attributes with high confidence (Brandao, 2019). Using the filtered datasets, the two annotators independently label the gender and age attributes for each image. For gender, we follow previous studies (Brandao, 2019) and consider only two labels: male and female. As for age, in line with the literature (Brandao, 2019; Kogure et al., 2022), we classify pedestrians into two labels: child and adult, based on their physical characteristics depicted in the images.
To ensure the reliability of the labeling procedure, both annotators independently label the age and gender attributes for each image. We use Cohen’s Kappa ($\kappa$) (Cohen, 1960), a widely-adopted metric for measuring inter-rater agreement, during the independent labeling process. The obtained $\kappa$ values for age and gender attributes in each of the four datasets are summarized in Table 3. According to the literature (Landis and Koch, 1977), a $\kappa$ value between 0.81 and 1 signifies almost perfect agreement, while a value between 0.61 and 0.8 indicates substantial agreement. In our labeling process, we achieve substantial agreement in gender labeling for the EuroCity-Day dataset and almost perfect agreement for all other tasks. This high level of agreement underscores the reliability of our labeling procedure (Wen et al., 2021; Chen et al., 2021). In cases where the two annotators encounter conflicts, an arbitrator is involved in the discussion to reach a consensus. After the labeling process, the summary of the number of labeled pedestrian instances for each dataset is presented in Table 4.
Table 3. Cohen’s $\kappa$ values for labeling age and gender.
Dataset $\kappa_{age}$ $\kappa_{gender}$
CityPersons 0.847 0.814
EuroCity-Day 0.925 0.800
EuroCity-Night 0.847 0.870
BDD100k 0.828 0.854
Table 4. Number of labeled pedestrian instances per dataset.
Gender Age Skin tone
Dataset Male Female Adult Child Light Dark
CityPersons 2,357 1,822 4,568 233 - -
EuroCity-Day 1,726 1,646 4,498 100 - -
EuroCity-Night 1,265 1,318 4,165 68 - -
BDD100k 3,457 2,479 6,293 190 2,724 789
Overall 8,805 7,265 19,524 591 2,724 789
3.3.3. Scenario Labeling
To deeply explore the fairness of pedestrian detectors across various driving scenarios (i.e., different brightness, contrast and weather conditions), we also need images labeled with scenario information. We therefore classify the images containing labeled pedestrians (5,933 out of the total 8,311 images in all datasets) into different scenarios, to enable our analysis in RQ2. An alternative approach to achieving this purpose is via generating images with different scenarios using existing test generation techniques from the autonomous driving literature. We do not choose this approach because generated images are not real images and can suffer from unnaturalness (Zhang et al., 2022).
Labeling brightness. Brightness represents the overall lightness or darkness of the image. To distinguish the brightness of the images, we categorize the datasets into day time and night time. As seen in Table 2, images in the CityPersons and EuroCity-Day datasets belong to day time, while images in the EuroCity-Night dataset belong to night time. For the BDD100k dataset, Wilson et al. (Wilson et al., 2019) have publicly provided the labels for day time and night time. We directly use the label information for study.
Labeling contrast. Contrast is the difference in brightness between objects in an image. To quantify the contrast of each image, we use the RMS contrast measurement (Peli, 1990), a standard measure in the computer vision literature (Bhuiyan and Khan, 2018; Jakhetiya et al., 2018; Pelli and Bex, 2013). To apply the RMS measurement, we first need to convert all images into the grayscale mode (Peli, 1990). Then, we calculate the RMS contrast value for each image based on the converted version. A higher RMS contrast value indicates a greater contrast. We partition the range of RMS contrast values into three classes, each class featuring equal value intervals. These classes correspond to distinct levels, labeled from level 1 to level 3, with higher levels indicating images with higher contrast. Then, we categorize the images into their respective contrast levels based on their RMS contrast values.
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2017). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because the two annotators and the arbitrator all cannot accurately distinguish the weather conditions.
To measure inter-rater agreement during manual labeling, we also use Cohen’s Kappa ($\kappa$). The $\kappa$ value is 0.813, indicating almost perfect agreement (Landis and Koch, 1977). This high level of agreement confirms the reliability of our labeling procedure. After scenario labeling, the summary of the number of images under different brightness, contrast, and weather conditions is presented in Table 5.
Table 5. Number of images in different brightness conditions, contrast levels, and weather conditions.
Brightness conditions Contrast levels Weather conditions
Day-time Night-time Level 1 Level 2 Level 3 Rainy Non-Rainy
4,409 1,524 1,163 3,933 837 277 3,800
3.4. Evaluation Metric
There have been well-established quantitative measures for group fairness in the literature. The most widely-adopted fairness measures include SPD (Statistical Parity Difference), EOD (Equal Opportunity Difference), and AOD (Average Odds Difference) (Chen et al., 2022, 2024; Zhang and Harman, 2021). SPD quantifies the difference in probabilities of favorable outcomes between unprivileged and privileged groups; EOD indicates the difference in the true-positive rate between unprivileged and privileged groups; AOD calculates the average difference between the false-positive rate and true-positive rate for unprivileged and privileged groups.
In the context of pedestrian detection, both SPD and EOD measure the disparity in proportions of successfully detected pedestrians between privileged and unprivileged groups. They also both express the difference in miss rates between privileged and unprivileged groups. Miss Rate (MR) is the most commonly-studied performance metric in pedestrian detection (Brandao, 2019; Kogure et al., 2022), which quantifies the proportion of undetected pedestrians. Formally, it is calculated as follows:
(1) $MR=1-\frac{TP}{TP+FN},$
where TP (true positives) refers to the number of successfully detected ground-truth bounding boxes, and FN (false negatives) denotes the number of undetected ground-truth bounding boxes. Pedestrian detectors generate bounding box locations and confidence scores for recognized “person” instances in images. To assess whether a given ground-truth bounding box is successfully detected, the standard method in the literature is to use the Intersection over Union (IoU) metric (Dollár et al., 2012). The IoU metric quantifies the degree of overlap between the ground-truth bounding box and the detected bounding box. If the IoU value is greater than 50%, the ground-truth bounding box is considered successfully detected (Dollár et al., 2012). Otherwise, it is classified as undetected.
The calculation of AOD requires precise false-positive information, referring to instances where members of the negative class (non-pedestrians) are incorrectly classified as the positive class (pedestrians). As described in Section 3.3.2, we adhere to the standard practice in the literature, where we focus on a “reasonable subset” of pedestrians within the images. This approach presents challenges in calculating precise false-positives for each group because the negative class (non-pedestrian) may also include instances that belong to the positive class (pedestrian). Therefore, we do not consider AOD in this study.
In summary, we use both SPD and EOD as fairness measures for evaluating pedestrian detectors. Since these two measures yield identical values for pedestrian detection, for the remainder of the paper, we present only EOD.
3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups), we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely used to analyze differences between proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if if the obtained $p$-value falls below a predetermined threshold (in our case, 0.05, a widely-accepted threshold in the fairness literature (Chen et al., 2022, 2023)). For instance, in evaluating whether there exists a difference between the miss rates for male ($MR_{male}$) and female ($MR_{female}$) individuals detected by a pedestrian detector, the null hypothesis assumes that $MR_{male}$ is equal to $MR_{female}$. If the resulting $p$-value is lower than 0.05, we reject the null hypothesis, indicating a significant difference between $MR_{male}$ and $MR_{female}$.
3.6. Experimental details
The experiments are implemented based on open-source frameworks of each pedestrian detector. For general object detectors, we select the pre-trained models with the highest accuracy from the the MMdetection model zoo  (MMDetection Developers, nd). For pedestrian-specific detectors, we employ pre-trained models available from their respective public repositories  (hasanirtiza, 2021; sxlpris, 2020; liuwei16, 2018).
To ensure the reliability of our results, all experiments are repeated ten times. The final results were derived by calculating the average across these ten iterations.
All experiments are performed on a platform equipped with 64GB RAM, 2.5GHz Intel Xeon (R) v3 Dual CPUs, and one NVIDIA GeForce RTX 2080 Ti GPU. YOLOX, RetinaNet, Faster RCNN, and Cascade RCNN are implemented using PyTorch 1.8.1 and Python 3.7 on Ubuntu 18.04 LTS, following the MMdetection configuration (Chen et al., 2019). CSP and MGAN use PyTorch 1.10.0 and Python 3.8 on Ubuntu 18.04 LTS, adhering to the Pedestron configuration (Hasan et al., 2022). ALFNet (Liu et al., 2018) and PRNet (Song et al., 2020) are implemented using Keras 2.0.6, Tensorflow 1.4.0, and Python 2.7 on Ubuntu 16.04 LTS.
4. Results
This section answers our RQs based on experimental results.
4.1. RQ1: Overall Fairness
RQ1 investigates the overall fairness of eight state-of-the-art pedestrian detectors regarding gender, age, and skin tone. First, for each detector, we compute the miss rate (MR) for different demographic groups and calculate EOD based on the MR results over all the datasets that we study. We also use the two-proportion z-test to determine the significance of any observed unfairness, as described in Section 3.5. Table 6 presents the results, with significant unfairness results highlighted in shading. In the following, we analyze the results for gender, age, and skin tone, respectively.
Table 6. (RQ1) Overall fairness in pedestrian detection across gender, age, and skin tone. Statistically significant biases, indicated by EOD (i.e., miss rate difference), are shaded. On average, detectors display comparable miss rates for female and male pedestrians, as well as for dark-skin and light-skin individuals. However, concerning age, significant bias is observed, as pedestrian detectors exhibit a 20.14% higher miss rate for children compared to adults.
Detectors MR Male MR Female EOD (Gender) MR Adult MR Child EOD (Age) MR Light-skin MR Dark-skin EOD (Skin tone)
YOLOX 9.78% 10.50% -0.72% 12.64% 42.47% -29.83% 5.21% 3.80% 1.41%
RetinaNet 10.72% 12.59% -1.87% 14.36% 44.33% -29.97% 8.33% 4.44% 3.90%
Faster RCNN 3.80% 4.13% -0.32% 5.24% 26.06% -20.82% 5.91% 3.30% 2.62%
Cascade RCNN 3.87% 4.16% -0.28% 5.12% 26.57% -21.44% 5.21% 3.04% 2.17%
ALFNet 30.86% 32.90% -2.04% 36.62% 53.47% -16.85% 42.55% 43.98% -1.43%
CSP 33.49% 35.25% -1.76% 37.74% 50.42% -12.68% 61.38% 64.77% -3.39%
MGAN 29.76% 30.97% -1.21% 33.58% 46.53% -12.95% 52.94% 54.88% -1.94%
PRNet 40.28% 41.62% -1.34% 44.69% 61.25% -16.56% 59.65% 59.44% 0.21%
Average 20.32% 21.52% -1.19% 23.75% 43.89% -20.14% 30.15% 29.71% 0.44%
[…]
Furthermore, we analyze the miss rate difference achieved by each pedestrian detector across the four datasets used in our study. Figure 2 illustrates the results. Significant gender biases (i.e., significant miss rate differences) are indicated by labeled miss rate values. We observe that in the CityPersons and EuroCity-Day datasets, containing only day-time data, only one detector in the EuroCity-Day datasets exhibits significant differences in miss rates between females and males. For the remaining results, there are no notable differences in miss rates between genders. However, results from the EuroCity-Night dataset present a contrasting observation, where seven of eight detectors exhibit a significantly higher miss rate for females, indicating bias in detecting females. In the BDD100k dataset that includes both day-time and night-time images, the miss rate difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
Figure 2. (RQ1) Miss rates of pedestrian detectors for females and males across datasets. Statistically significant gender biases are indicated by labeled miss rate values. In CityPersons and EuroCity-Day datasets with only day-time data, no significant gender biases are observed. However, in the EuroCity-Night dataset, seven out of eight detectors show significantly higher miss rates for females, revealing bias in female detection.
[…]
Furthermore, we illustrate the miss rate difference across four datasets in Figure 3, with significant age biases indicated by labeled miss rate values. We observe that the miss rate of children is consistently higher than that of adults across all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant miss rate difference (i.e., $p$-value ¡ 0.05). This indicates that the bias favoring adults is not specific to a particular dataset or detector, highlighting a strong unfairness between adults and children.
Figure 3. (RQ1) Miss rates of pedestrian detectors for children and adults across datasets. Statistically significant age biases are labeled with miss rate values. In 30 out of 32 scenarios (comprising four datasets and eight detectors), children have significantly higher miss rates than adults.
The age bias may be attributed to the inherent challenge of detecting small objects, owing to the limited information provided by small bounding boxes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To demonstrate this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 4, reveal a correspondence between the two distributions. Children as well as undetected pedestrians tend to have smaller bounding boxes.
Figure 4. (RQ1) Bounding box size distributions of adults and children (left) and bounding box size distributions of undetected and detected pedestrians (right). We observe that both children and undetected pedestrians tend to have smaller bounding boxes.
Skin tone. As displayed in Table 6, current pedestrian detectors exhibit minimal bias between light-skin and dark-skin individuals, with an average miss rate difference of just 0.44%. Specifically, among the four pedestrian-specific detectors, none show significant skin tone bias, whereas among the four general object detectors, three display significant skin tone bias, with miss rate differences ranging from 2.17% to 3.9%. These disparities are notably smaller than those observed in age bias.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ1: On average, the eight state-of-the-art pedestrian detectors under study exhibit no significant performance difference in across gender and skin tone, but notable bias across age. In particular, pedestrian detectors exhibit a 20.14% higher miss rate for children compared to adults.
4.2. RQ2: Fairness in Different Scenarios
RQ2 evaluates the fairness of state-of-the-art pedestrian detectors under different real-world autonomous driving scenarios.
4.2.1. RQ2.1: Different brightness conditions.
As introduced in Section 3.3.3, we consider two brightness conditions: day time and night time.
We first evaluate the overall miss rates of the eight pedestrian detectors under day time and night time. The results, presented in Table 7, show a noticeable increase in average miss rates during night time compared to day time for each demographic group. For example, the average miss rates for males and females at night are 33.20% and 36.57%, respectively, compared to 17.24% and 17.32% during the day. A similar pattern is observed for age and skin tone. This indicates that the transition from day time to night time influences the performance of pedestrian detectors, with statistically significant higher miss rates observed at night.
Then, we investigate whether the performance is equally decreased for different demographic groups. Specifically, we explore the fairness change from day time to night time. Table 7 shows the results, with statistically significant biases (i.e., miss rate differences) emphasized in shading.
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the nighttime condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences. This potentially worsens concerns regarding female safety during nighttime out, a prevalent societal issue (Wang et al., 2023).
For age, all detectors exhibit significant biases during both day time and night time. Moreover, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night with statistically significant. This suggests a higher probability of children being undetected during night time.
For skin tone, the miss rate difference between dark-skin and light-skin groups increases from day time to night time, with the average difference increasing from 0.15% at day time to 3.16% at night time. Nevertheless, the overall skin tone bias is not statistically significant during both day time and night time.
Table 7. (RQ2.1) Miss rates and EOD of each pedestrian detector under day time and night time. Statistically significant unfairness results are shaded. We find that reduced brightness conditions not only decrease the performance of pedestrian detectors but also exacerbate their biases. Notably, while pedestrian detectors generally do not exhibit significant gender bias during day time, biases against females become pronounced with six out of eight detectors showing significant biases during night time.
Gender
Detectors Day-time Night-time
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
YOLOX 9.73% 9.38% 0.35% 9.99% 14.53% -4.54%
RetinaNet 10.42% 11.79% -1.37% 11.99% 15.48% -3.49%
Faster RCNN 3.66% 3.59% 0.07% 4.41% 6.06% -1.66%
Cascade RCNN 3.58% 3.68% -0.10% 5.11% 5.87% -0.76%
ALFNet 24.52% 24.39% 0.13% 57.29% 63.42% -6.14%
CSP 26.76% 26.38% 0.38% 61.57% 67.09% -5.51%
MGAN 25.48% 26.29% -0.81% 47.59% 47.76% -0.17%
PRNet 33.73% 33.07% 0.66% 67.63% 72.33% -4.70%
Average 17.24% 17.32% -0.09% 33.20% 36.57% -3.37%
Age
Detectors Day-time Night-time
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
YOLOX 10.80% 40.77% -29.97% 17.81% 54.93% -37.12%
RetinaNet 12.73% 41.92% -29.19% 18.97% 61.97% -43.01%
Faster RCNN 4.50% 25.00% -20.50% 7.34% 33.80% -26.46%
Cascade RCNN 4.52% 25.58% -21.05% 6.81% 33.80% -26.99%
ALFNet 26.52% 49.42% -22.90% 65.10% 83.10% -18.00%
CSP 28.64% 46.92% -18.28% 63.42% 76.06% -12.64%
MGAN 27.87% 43.46% -15.59% 49.72% 69.01% -19.30%
PRNet 34.80% 56.92% -22.13% 72.60% 92.96% -20.36%
Average 18.80% 41.25% -22.45% 37.72% 63.20% -25.48%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors Day-time Night-time
MR LS MR DS EOD (Skin) MR LS MR DS EOD (Skin)
YOLOX 4.89% 4.03% 0.86% 7.06% 2.52% 4.53%
RetinaNet 7.44% 4.33% 3.11% 13.38% 5.04% 8.34%
Faster RCNN 5.36% 3.13% 2.23% 9.00% 4.20% 4.80%
Cascade RCNN 4.84% 2.99% 1.86% 7.30% 3.36% 3.94%
ALFNet 37.05% 38.66% -1.61% 73.48% 73.95% -0.47%
CSP 55.90% 59.85% -3.95% 92.21% 92.44% -0.22%
MGAN 47.12% 49.70% -2.58% 85.64% 84.03% 1.61%
PRNet 54.39% 54.63% -0.24% 89.29% 86.55% 2.74%
Average 27.31% 27.16% 0.15% 47.17% 44.01% 3.16%
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.1: Lower brightness conditions not only diminish the performance of pedestrian detectors but also exacerbate their bias. Particularly, during day time, pedestrian detectors generally do not exhibit significant gender bias, whereas six out of eight detectors demonstrate significant biases against females during night time.
4.2.2. RQ2.2: Different contrast levels.
Table 8. (RQ2.2) Miss rates and EOD of each pedestrian detector under different contrast levels. On average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
Gender
Detectors level 3 level 2 level 1
Male MR Female MR EOD Male MR Female MR EOD Male MR Female MR EOD
YOLOX 6.30% 5.21% 1.09% 9.42% 10.42% -1.00% 13.99% 14.71% -0.72%
RetinaNet 7.89% 9.46% -1.56% 10.47% 12.76% -2.29% 13.99% 14.24% -0.25%
Faster RCNN 3.83% 2.34% 1.49% 3.66% 4.16% -0.50% 4.34% 5.32% -0.98%
Cascade RCNN 3.43% 2.44% 0.98% 3.65% 4.26% -0.62% 5.12% 5.01% 0.11%
ALFNet 33.33% 31.67% 1.66% 30.28% 32.66% -2.38% 31.09% 34.74% -3.65%
CSP 40.11% 37.41% 2.70% 32.61% 34.80% -2.19% 31.54% 35.45% -3.90%
MGAN 38.52% 38.89% -0.38% 29.25% 31.23% -1.98% 24.61% 24.10% 0.51%
PRNet 48.41% 46.97% 1.43% 39.45% 41.52% -2.06% 36.92% 38.11% -1.19%
Average 22.73% 21.80% 0.93% 19.85% 21.48% -1.63% 20.20% 21.46% -1.26%
Age
Detectors level 3 level 2 level 1
Adult MR Child MR EOD Adult MR Child MR EOD Adult MR Child MR EOD
YOLOX 6.91% 32.89% -25.98% 12.05% 40.92% -28.87% 18.23% 53.23% -35.00%
RetinaNet 10.28% 36.84% -26.56% 13.85% 41.94% -28.10% 18.70% 56.45% -37.75%
Faster RCNN 4.00% 23.68% -19.69% 5.08% 26.09% -21.00% 6.56% 27.42% -20.86%
Cascade RCNN 3.33% 19.74% -16.41% 5.05% 25.58% -20.52% 6.48% 33.87% -27.39%
ALFNet 34.26% 52.63% -18.37% 35.42% 54.73% -19.32% 42.24% 50.00% -7.76%
CSP 40.72% 53.95% -13.23% 36.92% 52.17% -15.25% 38.72% 42.74% -4.02%
MGAN 40.38% $4$35% 33.20% 48.85% -15.65% 30.68% 40.32% -9.65%
PRNet 48.33% 68.42% -20.09% 43.46% 63.17% -19.71% 46.67% 50.81% -4.14%
Average 23.53% 41.61% -18.08% 23.13% 44.18% -21.05% 26.04% 44.35% -18.32%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors level 3 level 2 level 1
LS MR DS MR EOD LS MR DS MR EOD LS MR DS MR EOD
YOLOX 4.15% 4.91% -0.75% 5.75% 3.35% 2.40% 2.52% 2.17% 0.35%
RetinaNet 7.27% 5.28% 1.99% 8.60% 4.18% 4.41% 10.08% 2.17% 7.91%
Faster RCNN 5.04% 4.91% 0.14% 6.16% 2.51% 3.65% 6.72% 2.17% 4.55%
Cascade RCNN 4.01% 4.53% -0.52% 5.64% 2.09% 3.55% 5.04% 4.35% 0.69%
ALFNet 37.09% 36.60% 0.49% 42.62% 45.40% -2.78% 72.27% 71.74% 0.53%
CSP 53.26% 57.74% -4.47% 62.40% 65.48% -3.08% 90.76% 97.83% -7.07%
MGAN 49.26% 50.19% -0.93% 52.30% 54.81% -2.51% 84.03% 82.61% 1.42%
PRNet 54.15% 54.72% -0.56% 59.92% 59.21% 0.71% 86.55% 89.13% -2.58%
Average 26.78% 27.36% -0.58% 30.42% 29.63% 0.80% 44.75% 44.02% 0.73%
Following the roadmap outlined in RQ2.1, we begin by comparing the overall miss rates of pedestrian detectors across different contrast levels, as presented in Table 8. As explained in Section 3.3.3, we categorize driving scenarios into three contrast levels, with a higher level indicating greater contrast. However, we do not observe a consistent pattern in the results. Specifically, concerning light-skin and dark-skin pedestrians, we note a decrease in miss rates overall with increasing contrast. Nonetheless, this pattern is not observed across other demographic groups.
We then examine the shifts in fairness, as evidenced by the trends in EOD with increasing contrast, as outlined in Table 8. Our analysis reveals a consistent pattern: on average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
Specifically, for gender, detectors achieve -1.63% EOD under level 2 (most biased), -1.26% under level 1, and 0.93% under level 3 (fairest). For age, detectors achieve -21.05% EOD under level 2 (most biased), -18.32% under level 1, and 18.08% under level 3 (fairest). For skin tone, detectors achieve 0.80% EOD under level 2 (most biased), 0.73% under level 1, and -0.58% under level 3 (fairest).
This finding is further supported by the observation that under level 2, the highest number of detectors exhibit significant biases for gender, age, and skin tone, whereas under level 3, the fewest detectors show significant biases. Specifically, under level 2, five detectors demonstrate significant biases regarding gender, whereas under level 3, only one detector does so. Under level 2, four detectors show significant biases regarding skin tone, whereas under level 3, no detector exhibits significant biases.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.2: We classify driving scenarios into three contrast levels and observe that while there is not a clear pattern in the overall detection performance change with contrast variation, the highest contrast level consistently produces the fairest detection results across gender, age, and skin tone.
4.2.3. RQ2.3: Different weather conditions.
As decribed in Section 3.3.3, we consider two weather conditions: non-rainy and rainy.
Table 9 presents the miss rates of eight pedestrian detectors under the two conditions, with statistically significant unfairness results shaded. Overall, the miss rate for each demographic group increases in rainy weather conditions. This escalation could be attributed to droplets covering the camera and disrupting the detectors. A follow-up question is whether rainy weather fairly increase the miss rate for different demographic groups.
Table 9. (RQ2.3) Miss rates and EOD of each pedestrian detector under non-rainy and rainy weather conditions. Statistically significant unfairness results are shaded. We find that rainy weather does not largely impact fairness regarding gender, age, and skin tone.
Gender
Detectors Non-rainy weather Rainy weather
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
YOLOX 9.58% 8.95% 0.63% 8.33% 5.17% 3.16%
RetinaNet 10.62% 11.35% -0.74% 10.09% 8.05% 2.04%
Faster RCNN 3.68% 3.29% 0.39% 4.39% 2.87% 1.51%
Cascade RCNN 3.68% 3.48% 0.20% 4.39% 2.87% 1.51%
ALFNet 23.52% 22.92% 0.60% 42.54% 40.80% 1.74%
CSP 25.29% 24.35% 0.94% 50.44% 44.25% 6.19%
MGAN 24.63% 25.26% -0.62% 50.88% 50.57% 0.30%
PRNet 32.73% 31.64% 1.09% 51.75% 50.57% 1.18%
Average 16.72% 16.41% 0.31% 27.85% 25.65% 2.20%
Age
Detectors Non-rainy weather Rainy weather
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
YOLOX 10.45% 39.95% -29.50% 9.40% 30.00% -20.60%
RetinaNet 12.49% 40.65% -28.16% 13.61% 40.00% -26.39%
Faster RCNN 4.36% 23.09% -18.73% 4.86% 30.00% -25.14%
Cascade RCNN 4.53% 23.79% -19.25% 4.38% 25.00% -20.62%
ALFNet 25.00% 47.34% -22.34% 45.06% 75.00% -29.94%
CSP 26.86% 44.11% -17.25% 48.78% 65.00% -16.22%
MGAN 26.49% 41.11% -14.62% 53.00% 65.00% -12.00%
PRNet 33.19% 53.58% -20.39% 53.32% 75.00% -21.68%
Average 17.92% 39.20% -21.28% 29.05% 50.63% -21.57%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors Non-rainy weather Rainy weather
MR LS MR DS EOD (Skin) MR LS MR DS EOD (Skin)
YOLOX 4.19% 3.47% 0.72% 16.94% 13.51% 3.42%
RetinaNet 6.56% 3.81% 2.74% 21.77% 13.51% 8.26%
Faster RCNN 4.59% 2.60% 1.99% 15.32% 10.81% 4.51%
Cascade RCNN 3.93% 2.60% 1.33% 16.13% 8.11% 8.02%
ALFNet 34.90% 36.92% -2.02% 66.13% 64.86% 1.26%
CSP 53.05% 57.02% -3.97% 79.03% 89.19% -10.16%
MGAN 44.68% 47.83% -3.15% 72.58% 78.38% -5.80%
PRNet 51.99% 52.86% -0.87% 77.42% 81.08% -3.66%
Average 25.49% 25.89% -0.40% 45.67% 44.93% 0.73%
From Table 9, we observe that rainy weather may potentially mitigate bias in pedestrian detectors. Specifically, under non-rainy conditions, all eight detectors exhibit significant bias against children, while under rainy conditions, three out of four pedestrian-specific detectors no longer show significant age bias any more. Similarly, under non-rainy conditions, two of eight detectors display significant bias against skin tone, whereas under rainy conditions, none do. However, the improvement in fairness due to rainy weather is marginal. The EOD difference between rainy and non-rainy conditions is 1.89% for gender, -0.29% for age, and 1.13% for skin tone.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.3: Rainy weather conditions decrease overall detection performance to a large extent but have a subtle impact on the fairness of pedestrian detectors. Specifically, three out of four pedestrian-specific detectors no longer exhibit significant age bias under rainy weather. However, the improvement in fairness is marginal.
5. Discussion
5.1. Fairness-performance Trade-off
It is widely acknowledged that fairness improvement usually comes at the cost of machine learning performance (e.g., accuracy) (Ge et al., 2022; Wick et al., 2019; Chen et al., 2023; Hort et al., 2021). Therefore, developers need to grapple with the challenge of optimizing ML performance without compromising fairness, encapsulating this tension as the “fairness-performance trade-off.”
Nonetheless, we have observed that pedestrian detectors exhibiting greater fairness (i.e., lower absolute EOD values) can achieve superior performance (i.e., lower overall miss rates) under certain environmental conditions. For instance, in the results of RQ2.1, we find that during daytime compared to nighttime, the eight pedestrian detectors achieve better detection results while simultaneously displaying reduced EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2 reveal that the highest level of contrast consistently leads to improved fairness regarding gender, age, and skin tone, while also demonstrating the best overall detection performance in specific cases. These findings offer a positive counterpoint to previous fairness-performance trade-off trade-off theory, suggesting that we can enhance both detection performance and fairness by adjusting the brightness and contrast of captured images.
5.2. Implications
5.2.1. Implication for researchers
Our empirical results offer valuable research opportunities for researchers. 1) Fairness improvement of autonomous driving systems via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific post-processing image editing techniques to adjust contrast and brightness of captured input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Real-time adaptive fairness improvement of autonomous driving systems. Adapting software systems to increasingly dynamic environments poses significant challenges in software development, an area that has long been a important task in SE research (Qureshi and Perini, 2009). Our empirical findings contribute to this ongoing task by introducing a new challenge: designing real-time adaptive fairness improvement methods for autonomous driving systems. Specifically, our research reveals that autonomous driving systems exhibit greater biases under specific driving scenarios influenced by changes in brightness and contrast. Given the common occurrence of such changes in real-world driving conditions, there is an urgent need to develop specific methods to ensure the consistent fairness of autonomous driving systems in response to these environmental dynamics. 3) Fairness improvement of autonomous driving systems via multi-objective optimization. Given our finding that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. 4) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 5) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage fellow researchers to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
5.2.2. Implication for developers
Fairness is a critical non-functional requirement for software applications, but our study demonstrates the existence of significant bias in state-of-the-art pedestrian detectors. It is imperative for developers to prioritize their efforts in this domain. Addressing these concerns goes beyond enhancing the quality of autonomous driving systems; it also serves to safeguard autonomous driving companies from ethical, reputational, financial, and legal repercussions that may arise in the event of violations of anti-discrimination laws. Moreover, it is crucial for developers of autonomous driving systems to consider the influence of various environmental factors on fairness during the development process.
5.2.3. Implication for policy makers
Governments have a role in raising awareness about the potential biases in autonomous driving systems. 1) Regulatory measures. Autonomous driving systems play a crucial role in ensuring human safety. Our results reveal that fairness issues exist in modern pedestrian detection models. Therefore, it is essential for policy makers to enact regulations and standards that safeguard the rights of all individuals and address these concerns appropriately. 2) Enhanced protection for vulnerable pedestrians. Policies should also aim to address safety issues for vulnerable pedestrians such as children and females, who are, as revealed by this study, more likely to be overlooked by the current detection models.
5.3. Threats to Validity
Manual labeling. The labeling process involves possible subjectivity, posing threats to the validity of the analysis. To mitigate this threat, two annotators and an arbitrator are involved in the labeling process: first each image is independently labeled by two annotators, and in cases where conflicts in labeling arise, we seek the expertise of an arbitrator to resolve such discrepancies and arrive at a consensus. The inter-rater agreement between the two annotators is high, which demonstrates the reliability of the labeling schema and procedure adopted herein.
Selection of pedestrian detectors. Our study is based on eight pedestrian detectors, which may lead to possible selection bias in this study. To mitigate this threat, we select representative pedestrian detectors based on two considerations. On one hand, the selected detectors are widely-studied in pedestrian detection and autonomous driving literature (Cao et al., 2020; Zhang et al., 2016, 2017; Yu et al., 2020). On the other hand, the selected models cover the two typical types of pedestrian detection methods (i.e., one-stage detection and two-stage detection), ensuring a comprehensive representation of the techniques used in the field.
Selection of datasets. The choice of datasets presents a potential threat to validity. To address this concern, we use four real-world datasets extensively explored in autonomous driving research, including a total of 8,311 images. We augment the datasets with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. The extensive scale of data ensures the reliability of our results.
Selection of evaluation measures. The fairness measures that we employ may introduce potential limitations. To mitigate this concern, we consider three commonly-used fairness measures: SPD, EOD, and AOD. Upon careful discussion, we conclude that AOD may not be suitable for our study (Section 3.4), leading us to focus on SPD and EOD. We demonstrate that SPD and EOD lead to equivalent observations for pedestrian detection. Ultimately, we opt to use EOD (i.e., SPD) for our analysis, as it involves the comparison of the miss rates of different demographic groups. This ensures a consistent evaluation with the current pedestrian detection research, as miss rate is the most widely-adopted metric for measuring the performance of pedestrian detectors in the literature (Brandao, 2019; Kogure et al., 2022).
6. Conclusion
This paper presents the first comprehensive study on fairness of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors regarding gender, age, and skin tone. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. Additionally, pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Regarding skin tone, we observe balanced detection performance for light-skin and dark-skin individuals overall. As part of our contribution, we publicly release large-scale real-world pedestrian detection datasets with gender, age, and skin tone labels. These datasets aim to facilitate future fairness research in autonomous driving. The insights gained in this study can pave the way for more fair and unbiased autonomous driving systems in the future.
7. Data Availability
Our GitHub repository (git, 2024) contains datasets, sensitive attribute labels, scripts, and results of this work to facilitate replication and extension.